94g NaHCO_(3)*(1 mole NaHCO_(3))/(84g NaHCO_(3))*(1 mole NaCl)/(1 mole NaHCO_(3))*(58.5g NaCl)/(1 mole NaCl)
1
Expert's answer
2020-11-02T08:53:48-0500

"\begin{aligned} 
&94g NaHCO_3 × \dfrac{1\ mole \ NaHCO_3}{84g\ NaHCO_3}×\\
\\
&\dfrac{1\ mole \ NaCl}{1\ mole\ NaHCO_3} × \dfrac{58.5g \ NaCl}{1\ mole\ NaCl} =
\end{aligned}"


Re-arranging the numerators and denominations, so that"NaHCO_3" and "NaCl" are grouped together, and moles and masses are also grouped together, we have;

"\begin{aligned} 
&\dfrac{1\ mole\ of\ NaHCO_3}{1\ mole\ NaHCO_3} × \dfrac{94g NaHCO_3 }{84g\ NaHCO_3}×\\
\\
&\dfrac{1\ mole \ NaCl}{1\ mole\ NaCl} × 58.5g \ NaCl= 
\end{aligned}"


"\begin{aligned}
1 ×1.12 × 1 ×58.5g \ NaCl= 
\end{aligned}"

"65.5g\ of\ NaCl"